Can you help me solve this problem?
I am suppose to solve 2x2 by elimination method.
2p-5q = -33
6p-7q =39

Thank you,
Yolanda

Answer by checkley71(8403)

MULTIPLY THE FIRST EQUATION BY -3 & ADD.
-6P+15Q=99
6P-7Q=39
--------------------------
8Q=138
Q=138/8
Q=17.25 ANSWER.
6P-7*17.25=39
6P-120.75=39
6P=39+120.75
6P=159.75
P=159.75/6
P=26.625 ANSWER.
PROOF:
2*26.625-5*17.25=-33
53.25-86.25=-33
-33=-33